Suppose a soap-manufacturing production process is described by the following equation:

Y = a + b log K + с log L

Where,

Y= Output (number of soaps produced) K=Capital

L=Labor

a, b and c are constants

Suppose 0<a<1, 0< b<1 and 0<c<1

  1. Find the Marginal Product of Labor (MPL) and Marginal Product of Capital (MPK) in the production of soap
  2. Is MPL diminishing, increasing or constant as L increases?
  3. Is MPK diminishing, increasing or constant as K increases? 
1
Expert's answer
2021-10-10T16:35:24-0400

1. The marginal product of labor (MPL) is calculated as follows:

"MP_L =\frac{ ∂Y}{∂L}\\=\frac{∂(a + b\space log\space K + c\space logL)}{∂L}\\=\frac{c}{L}"

The marginal product of labor (MPK) is calculated as follows:

"MP_K =\frac{ ∂Y}{∂K}\\=\frac{∂(a + b\space log\space K + c\space logL)}{∂L}\\=\frac{b}{K}"


2. The marginal product of labor is:

"MP_L =\frac{ c}{L}"

The labor is in denominator. If L increases then MPL decreases. Thus, it is diminishing.

3. The marginal product of capital is:

"MP_k =\frac{ b}{K}"

The capital is in denominator. If K increases then MPK decreases. Thus, it is diminishing.